OpenStudy (anonymous):

3m+5=8m-105 solve plz

OpenStudy (jadeishere):

3m + 5 = 8m - 105 3m -(8m) + 5 = 8m -(8m) -105 -5m + 5 = -105 -5m = -110 m = -110/-5 m = 22

OpenStudy (anonymous):

RULE to use here: subtract, add, multiply and divide to get numbers on one side and variables on the other side

OpenStudy (anonymous):

RULE #2:: divide each side by the coefficient of the variable to completely isolate the variable (f.e. 2m=4--->divide each side by 2 to isolate m)